The molecular mass of CuSO4.5H2O is:

(ATOMIC MASS OF CU = 63.5, H = 1, O = 16, S = 32)
A

233.5 gm/mole

B

249.5 gm/mole

C

251.5 gm/mole

D

247.5 gm/mole

 

Grade:8

1 Answers

Omveer Singh
3 years ago

Molecular mass of a compound is the sum of atomic masses of each element present in the molecule. 

Molecular mass = atomic mass of Cu + atomic mass of S + (9 × atomic mass of O) + (10 × atomic mass of H=63.5+32+9×16+10×1=249.5 gm/mole